OpenStudy (anonymous):

Can someone help me with 2 more questions Medal rewarded for correct answers @BioEpic @iGreen @Haseeb96 @ashsek MEDALS ONLY REWARDED FOR CORRECT ANSWERS (QUSTIONS BELOW)

OpenStudy (anonymous):

@Haseeb96 @BioEpic @iGreen @ashsek

OpenStudy (anonymous):

I got 1792

OpenStudy (igreen):

\(V = lwh\) \(8 \cdot 14 \cdot 16 = 1792~\huge\color{lime} \checkmark\)

OpenStudy (anonymous):

Okay @iGreen one more and you will get a medal

OpenStudy (anonymous):

@iGreen and @Haseeb96

OpenStudy (igreen):

Formula: \(V = \dfrac{1}{2} l w h\)

OpenStudy (haseeb96):

Volume = 1/2 L x W x H put the values in this formula and get the answer

OpenStudy (anonymous):

Is it WxLxH

OpenStudy (igreen):

\(V = \dfrac{1}{2} l w h\) \(V = \dfrac{1}{2} (30) (12) (8)\) \(V = (15)(12)(8)\) You can do the rest.

OpenStudy (anonymous):

2880

OpenStudy (igreen):

No, check again.

OpenStudy (anonymous):

@Haseeb96 2880

OpenStudy (igreen):

Oh, well 30 * 12 * 8 is 2880, but you have to multiply 2880 by 1/2, can you do that?

OpenStudy (anonymous):

1440

OpenStudy (igreen):

\(2880 \cdot \dfrac{1}{2} = 1440~\huge\color{lime} \checkmark\)

OpenStudy (haseeb96):

yes this is correct answer Good job again girl

OpenStudy (anonymous):

U each got a medal good job

OpenStudy (haseeb96):

haha :) would u help me in writing a good essay @starrhazel123

OpenStudy (anonymous):

yes

OpenStudy (haseeb96):

because it is not my language so i dont write in good manner

OpenStudy (haseeb96):

so would u tell me from what day should we start it?

OpenStudy (anonymous):

umm whenever how about tommrow around lunch or tonight at 10:30 pm

OpenStudy (anonymous):

or Monday ill text you k

OpenStudy (haseeb96):

okay i will write an essay and u will tell me how can u improve it and what are my mistakes in that essay .

OpenStudy (anonymous):

okay tommrow around 9:30 k

OpenStudy (haseeb96):

actually i m not from USA i m from Asia so here the timing is different so tomorrow 9:30 am or pm of your country

OpenStudy (anonymous):

Am in my country

OpenStudy (haseeb96):

okay what is the time there right now/

OpenStudy (anonymous):

11:17 am

OpenStudy (haseeb96):

okay here 9:16 pm

OpenStudy (haseeb96):

do u have cell phone ?

OpenStudy (anonymous):

Okay I will meet you at launguege arts and open up a question and tag you in it and then we will disscuss

OpenStudy (haseeb96):

okay thanks also in advance :) u are the great girl